A train traveling at 30mph leaves for a certain town.Two hours later a bus traveling 40mph leaves for the same town and arrives at the same time as the train.If both the train and the bus traveled in a straight line, how far is the town from where they started
Answer by scott8148(6628)

t = d / r

(d / 30) + 2 = d / 40